OpenStudy (anonymous):

Find all the zeros of the equation. -4x4 - 44x2 + 3600 = 0

OpenStudy (anonymous):

-4x^4 - 44x^2

OpenStudy (anonymous):

\[-4 x^4-44 x^2+3600=-4 (x-5) (x+5) \left(x^2+36\right) \]